OpenStudy (rainbowdashrocks):

Estimate by first rounding each number to the nearest 1/2 . -4 1/6 * (-1 14/15 A.-12 B.-8 C.8 D.12 WILL FAN AND MEDAL PLZ HELP

OpenStudy (rainbowdashrocks):

@Thadds2003

OpenStudy (anonymous):

B I think

zepdrix (zepdrix):

\[\large\rm -4\frac16\times\left(-1\frac{14}{15}\right)\]These are mixed numbers? Like this?

OpenStudy (rainbowdashrocks):

yas

OpenStudy (anonymous):

Wait, I didnt see the first - sign. Its C

zepdrix (zepdrix):

Might be easier to deal with if you convert to improper fractions :) Remember how to do that?

OpenStudy (rainbowdashrocks):

yes

OpenStudy (rainbowdashrocks):

i think its c

OpenStudy (rainbowdashrocks):

am i right

zepdrix (zepdrix):

Oh you were supposed to do rounding FIRST, ah i misread the question :( So -4 and 1/6 rounds up to -4. And -1 and 14/15 rounds down to -2. -4*-2 = 8 Yes, good job :)